# BEATS: Optimizing LLM Mathematical Capabilities with BackVerify and Adaptive Disambiguate based Efficient Tree Search

### v1

@@ 4.1 Experiment Settings
We conduct experiments on five authority mathematical reasoning datasets: (1) GSM8K: The GSM8K dataset consists of 1,319 test samples. It is widely used for arithmetic problem-solving tasks and is designed to evaluate models’ performance on grade-school-level math problems. (2) MATH: The MATH dataset contains 5,000 test samples, drawn from competition-style problems. It covers a wide range of topics, including algebra, calculus, combinatorics, and geometry. (3) SVAMP: The SVAMP dataset contains 1,000 math word problems, each with at most two mathematical expressions and one unknown variable. (4) SimulEq: The SimulEq dataset includes 514 test samples. It centers on solving equations, with an emphasis on algebraic manipulation and logical reasoning. (5) NumGLUE: The NumGLUE dataset includes 1042 test problems, comprising 8 distinct tasks involving a range of numerical reasoning challenges, including arithmetic and quantitative reasoning within common sense, domain-specific contexts, reading comprehension, and natural language inference.
@@
To demonstrate the effectiveness of our approach, we conducted experiments using three different models: LLaMA3-8B-Instruct, Yi-1.5-6B-Chat and Qwen2-7B-Instruct. The main experimental results are presented in Figure 1 and Figure 2. The detaild analysis can be found in Section 4.2.
@@
We consider three types of baseline models: (1) Zero-Shot Models, which include Zero-Shot CoT and self-consistency based approache which first generate a set of candidate answers through multiple sampling and determine the final answer by majority voting. (2) Supervised Fine-Tuning (SFT) Models, which encompass WizardMath, MuggleMath, MetaMath, and LEMA-LLaMA. (3) Search Algorithm-Based Models, including ToT Yao et al. (2024), RAP Hao et al. (2023), ReST-MCTS* Zhang et al. (2024), LiteSearch Wang et al. (2024), and Llama-2+M*.
@@
In our experimental setup, we configured the tree depth to 7, with the disambiguation step allowed only as a direct successor to the root node. Node expansion was conducted using the vLLM framework, with the following parameters: temperature is 0.8, top_p is 0.9, and max_tokens is 2048. During the BackVerify stage, we employed Qwen2-7B-Instruct as the discriminator. For answer verification, we utilized the same framework as MAmmoth. All experiments were conducted on NVIDIA H800 GPU machines.
@@ 4.3 Ablation Study
Analysis of Disambiguation
Figure 3: From this figure, we observe that models are prone to errors when using majority voting but can achieve the correct answer through back verification.
Figure 4: From this figure, we observe that some questions may contain ambiguity, which can be resolved by using the disambiguation module to generate a clarified version of the question.
To evaluate the impact of the disambiguation process, we performed several comparative experiments using the MATH and GSM8K datasets with both LLaMA3 and Qwen2 models. As shown in Table 3, removing the disambiguation component in BEATS led to a noticeable decline in accuracy across all experiments, thereby confirming the importance of the disambiguation process. Additionally, we examined the effectiveness of disambiguation through case studies. In Figure 4, the clarified question offers the following advantages: 1) The original phrasing of "3 sprints 3 times a week" is ambiguous, as it could suggest that James runs three sprints three times a week, or that each session consists of three sets of three sprints. In contrast, the clarified question explicitly states that James runs three sprints per session and completes these sessions three times per week, reducing potential misinterpretation. 2) The clarified question concisely outlines the key details, "3 sprints of 60 meters each, 3 times a week", in a structured format that improves logical flow and understanding.
@@
MATH LLaMA3 w.o. disambiguation 23.2
BEATS 42.93
Qwen2 w.o. disambiguation 51.88
BEATS 61.52
GSM8K LLaMA3 w.o. disambiguation 74.83
BEATS 88.48
Qwen2 w.o. disambiguation 76.88
BEATS 83.02
Analysis of Back Verification
@@ 5 Conclusion
In this paper, we introduced BEATS, a new method designed to enhance the mathematical problem-solving capabilities of LLMs. By addressing critical challenges such as suboptimal prompts, ineffective verification methods, and high computational costs, our approach offers a significant improvement in performance. The meticulously crafted prompts facilitate step-by-step reasoning, reducing ambiguities in problem statements and enabling the model to generate accurate answers. Our innovative back-verification technique enhances the reliability of results by ensuring that answers are thoroughly validated. Additionally, the pruning tree search strategy allows for controlled inference time while maintaining state-of-the-art performance. Through extensive experimentation, we demonstrated that BEATS notably outperforms existing methods, marking a substantial step forward in the quest for improved mathematical reasoning in LLMs. Future work will explore further refinements to these techniques and their applicability across a wider range of complex problem domains.
Appendix A Prompt
Figure 5: Check wrong answer derived by majority voting.

### v2

@@ 4.1 Experiment Settings
We conduct experiments on five authoritative mathematical reasoning datasets: (1) GSM8K: The GSM8K dataset consists of 1,319 test samples and is widely used for arithmetic problem-solving tasks, designed to evaluate models’ performance on grade-school-level math problems. (2) MATH: The MATH dataset contains 5,000 test samples drawn from competition-style problems, covering a wide range of topics, including algebra, calculus, combinatorics, and geometry. (3) SVAMP: The SVAMP dataset comprises 1,000 math word problems, each involving at most two mathematical expressions and one unknown variable. (4) SimulEq: The SimulEq dataset includes 514 test samples focused on solving equations, with an emphasis on algebraic manipulation and logical reasoning. (5) NumGLUE: The NumGLUE dataset includes 1,042 test problems encompassing 8 distinct tasks that involve various numerical reasoning challenges, such as arithmetic, quantitative reasoning in commonsense and domain-specific contexts, reading comprehension, and natural language inference.
@@
To evaluate the effectiveness of our approach, we conducted experiments using three state-of-the-art (SOTA) models: LLaMA3-8B-Instruct, Yi-1.5-6B-Chat, and Qwen2-7B-Instruct. The primary experimental results are presented in Table 1 and Table 2, while a detailed analysis is provided in Section 4.2.
@@
We consider three types of baseline models: (1) Zero-Shot Models, which include Zero-Shot CoT and a hard-voting approach that first generates a set of candidate answers through multiple sampling and then determines the final answer by majority voting. (2) Supervised Fine-Tuning Models, encompassing WizardMath (Luo et al., 2023), MuggleMath (Li et al., 2024b), MetaMath (Yu et al., 2023), LEMA-LLaMA (An et al., 2023), Code-Llama (Roziere et al., 2023), Platypus (Lee et al., 2023), MAmmoTH (Yue et al., 2023), Galactica (Taylor et al., 2022), Tulu (Wang et al., 2023b), and Guanaco (Dettmers et al., 2023). (3) Search Algorithm-Based Models, including ToT (Yao et al., 2024), RAP (Hao et al., 2023), ReST-MCTS* (Zhang et al., 2024), LiteSearch (Wang et al., 2024), and Llama-2+M* (Kang et al., 2024).
@@
In our experimental setup, we configured the tree depth to 7, with the disambiguation step allowed only as a direct successor to the root node. Node expansion was performed using the vLLM framework with the following parameters: temperature set to 0.8, top_p set to 0.9, and max_tokens set to 2048. During the BackVerify stage, Qwen2-7B-Instruct was used as the discriminator. For answer verification, we employed the same framework as MAmmoth. All experiments were conducted on a machine running Ubuntu 22.04, equipped with 8 NVIDIA H100 GPUs, a 120-core CPU, and 960 GB of memory.
@@ 4.2 Main Experiment
Figure 3: From this figure, we observe that models are more likely to deduce errors when using majority voting but can achieve the correct answer through back verification.
@@ 4.3 Ablation Study
Remove the Disambiguation Module
Figure 4: From this figure, we observe that some questions may contain ambiguity, which can be resolved by using the disambiguation operation to generate a clarified version of the question.
To assess the impact of the disambiguation process, we conducted a series of comparative experiments using the MATH and GSM8K datasets with both the LLaMA3-8b-Instruct and Qwen2-7b-Instruct models. As shown in Table 3, removing the disambiguation component in BEATS resulted in a significant decrease in accuracy across all experiments, highlighting the critical role of the disambiguation process. Additionally, we evaluated the effectiveness of disambiguation through case studies. In Figure 4, the clarified question offers the following advantages: (1) The original phrasing, "3 sprints 3 times a week", is ambiguous, as it could imply that James runs three sprints three times a week or that each session consists of three sets of three sprints. In contrast, the clarified question explicitly states that James runs three sprints per session and completes these sessions three times per week, thereby minimizing potential misinterpretation. (2) The clarified question concisely presents the key details, "3 sprints of 60 meters each, 3 times a week", in a structured format that enhances logical flow and comprehension.
@@
MATH LLaMA3 BEATS 42.93
w.o. disambiguation 35.80 $\downarrow$ 7.13
Qwen2 BEATS 61.52
w.o. disambiguation 51.88 $\downarrow$ 9.64
GSM8K LLaMA3 BEATS 88.48
w.o. disambiguation 74.83 $\downarrow$ 13.65
Qwen2 BEATS 83.02
w.o. disambiguation 76.88 $\downarrow$ 6.14
Remove the Back Verification Module
@@ 5 Conclusion
In this paper, we introduced BEATS, a new method designed to enhance the mathematical problem-solving capabilities of LLMs. By addressing critical challenges such as suboptimal prompts, ineffective verification methods, and high computational costs, our approach offers a significant improvement in performance. The meticulously crafted prompts facilitate step-by-step reasoning, reducing ambiguities in problem statements and enabling the model to generate accurate answers. Our innovative back-verification technique enhances the reliability of results by ensuring that answers are thoroughly validated. Additionally, the pruning tree search strategy allows for controlled inference time while maintaining state-of-the-art performance. Through extensive experimentation, we demonstrated that BEATS notably outperforms existing methods, marking a solid foundation for advancing mathematical reasoning in LLMs. This work represents an excellent starting point, paving the way for future research to explore more effective verification methods and their applicability across a broader spectrum of complex problem domains.
Appendix A Search Cost at Inference Phase
Figure 5: Average tokens needed for solving different problems.
BEATS significantly improves the model’s mathematical capabilities through designed pruning search algorithm, which processes multi-turn question inference. Figure 5 presents a comparison of the average number of tokens generated by different models—LLaMA3, Qwen2, and Yi-1.5—across five mathematical benchmarks: MATH, GSM8K, SVAMP, SimulEq, and NumGLUE. As shown in the figure, LLaMA3 consistently produces the highest number of tokens across all benchmarks, with a particularly large margin in the MATH dataset, where it exceeds 5,000 tokens on average. In contrast, Qwen2 and Yi-1.5 generate fewer tokens, with Yi-1.5 often producing the least across most datasets. This suggests that LLaMA3 might engage in more extensive reasoning processes but at the cost of higher computation, while Qwen2 and Yi-1.5 strike a balance between efficiency and performance.
Appendix B Candidate Answer Distribution
Figure 6: Candidate answer set size.
Figure 6 illustrates the distribution of candidate answer set sizes for individual test samples across five mathematical benchmarks (MATH, GSM8K, SVAMP, SimulEq, and NumGLUE) for three models: LLaMA3, Qwen2, and Yi-1.5. As shown in the figure, most test samples for all models tend to have larger candidate sets, with a clear peak at 12 candidates across all benchmarks. LLaMA3 consistently demonstrates larger candidate sets compared to Qwen2 and Yi-1.5, particularly in the MATH and GSM8K benchmarks, where the size of candidate sets reaches up to 12 for a substantial number of cases.
Appendix C Prompts
Inspired by Li et al. (2024a), we utilized the prompts shown in Figure 7 to implement the BEATS algorithm.
Figure 7: Prompts used in BEATS.